OpenStudy (anonymous):

Simplify. 3sin(pi/12)cos(pi/12). I come up with (squareroot2)/2. How come this is incorrect?

OpenStudy (anonymous):

are you sure is pi/12?

OpenStudy (anonymous):

ok nvm got it one minute

OpenStudy (anonymous):

we need to use the identity that sin2x = 2sinxcosx alright so in this case we dont have a 2 there so we can right 1/2sin2x=sinxcosx alright multiplied by 3 on both sides its 3/2sin2(pi/12) = sin(pi/12)cos(pi/12 = 3/2sin(pi/6) we know sin (pi/6) = 1/2 by the special triangle so now we got (3/2)*(1/2) = 3/4 :)

OpenStudy (anonymous):

I see! That makes sense. I did my steps all wrong which explains why it is incorrect. Lol. Thank you so much :)